# DMUS-MP-25/04

### v1

@@ Abstract
The massless $S$-matrix of the pure RR $AdS_{3}\times S^{3}\times T^{4}$ theory is very similar but not quite the same as Fendley-Intriligator’s ${\cal{N}}=2$ $S$-matrix, in turns related to Sine-Gordon taken at a special coupling. In this short note we review the reason why supersymmetry emerges but with a fractional statistics of the particles. We then use this to obtain a very simple interpolating $S$-matrix between massless $AdS_{3}$ and Sine-Gordon, and further find two interpolating $S$-matrices between these and the mixed-flux relativistic $S$-matrix, solving the Yang-Baxter equation all the way in-between. They are all relativistic invariant and braiding unitary, can be written in terms of particles with fractional statistics, and they might serve the purpose to embed into a parent integrable system.
1 Introduction and Conclusions
It is undeniable that integrability of strings propagating on $AdS_{3}\times S^{3}\times T^{4}$ [1] (for reviews see [2, 3, 4]) continues to offer opportunities for surprises, keeping on with the tradition inaugurated by $AdS_{5}$ and $AdS_{4}$ [5]. The exact $S$-matrix theory is one of the richest available [6, 7, 8, 9, 10, 11] and the presence of massless modes [12, 13, 14], see also [15] and [16, 17], keeps pushing the technology further and further. The Quantum Spectral Curve (QSC) [18], see also [19], is undergoing an expansion of its methods to accommodate this model [20]. Meanwhile [21], see also [22, 23, 24], have considerably revitalised the field with new algebraic data and have created an exciting race with significant progress on all sides [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35].
@@ 1 Introduction and Conclusions
$E=\sqrt{\Big{(}m+\frac{k}{2\pi}p\Big{)}^{2}+4h^{2}\sin^{2}\frac{p}{2}}.$ (1.1)
@@
Notably $k\in\mathbbmss${Z} stands for the WZW level and $m$ for is the mass of the particle. An important wealth of work on this and similar deformations has now been done [42]. As it is customary, one calls pure RR the $k=0$ case (and mixed-flux otherwise).
@@
There is a special subarea inside this field which is evolving into its own little garden, and that is the relativistic sector. This goes from the BMN limit being non-trivial for same-chirality massless scattering [14], where it describes a CFT á lá Zamolodchikov [43], to the entire massless sector being of difference form [34, 35, 13, 21], to the mixed-flux scattering being amenable to a variety of massive relativistic limits [35, 23, 24], all pointing toward famous integrable $S$-matrices of the past - but with differences. The observation [14] is that the massless $S$-matrix of the pure RR theory is very similar but not quite the same as Fendley-Intriligator ${\cal{N}}=2$ $S$-matrix [44, 45, 46], in turns related to Sine-Gordon taken at a special coupling. Considerations on these connections in the case of mixed-flux are also expanded upon in [23], and in particular in [24], where the emphasis is on the differences in the global structure of the supersymmetry algebra.
@@
The focus of this paper is instead the difference in the nature itself of the supersymmetry, in particular the appearance of fractional statistics. We review the detailed reason why supersymmetry accidental pops out but is not exactly that which one has in pure RR $AdS_{3}$, due to a different assignment of fermionic number of the particles. We use the coproduct of the underlying quantum affine symmetry of Sine-Gordon to show this, following the discoverer of this phenomenon Bernard and LeClair [47], see also section 2 of [44] and [48]. We use this to obtain a very simple interpolating $S$-matrix between massless $AdS_{3}$ and Sine-Gordon, which could be useful.
@@
Then we felt inspired and deformed this interpolating $S$-matrix to include something like a non special coupling, basically relaxing the special coupling and letting it be generic. The hope was to obtain an interpolation to the (or one of the) mixed flux $S$-matrix (-ces). It did not exactly work like that, but we managed to find a couple of interpolations which do the job anyway, and solve the Yang-Baxter equation (YBE) all the way in-between. They are all relativistic invariant, and satisfy phyical unitarity for a choice of parameters. This means that we have found a few (perhaps new) relativistic integrable systems ($S$-matrices solving YBE) interpolating Fendley-Intriligator, massless $AdS_{3}$ and mixed flux relativistic. They are all regular, namely $S(\theta)\to\mathbbmss{1}\otimes\mathbbmss{1}$ as $\theta\to 0$, which means that they should sit within the classification of [49], possibly section 4.3.1 modulo twists [^1] They are also all of free-fermion type, which is consistent with having all the AdS/CFT $S$-matrix of free-fermion type - and also the Sine-Gordon $S$-matrix at special coupling $\xi\to 2\pi$, since Sine-Gordon satisfies $\mbox{const.}=4\cos^{2}\frac{\pi^{2}}{\xi}$ in formula (2.2) of [30]. They all have a semblance of $\mathfrak{sl}_{2}$-type quantum symmetry, they are braiding unitarity in deference to their quantum group origin, and they can be rephrased in terms of particles with fractional statistics [^2] Statistics in $1+1$ dimension is a fascinating subject, see for instance [50, 33].. Inspired by [33], we highlight this fact by calling them all fractional $S$-matrices - what seems to be a ubiquitous feature at this point [45].
Ultimately it is quite expected that a minimum continuation/deformation of sorts would fill up the gaps between these models. We believe however that it serves a useful purpose to explicitly embed all the relativistic $S$-matrices appearing from the study of string theory in $AdS_{3}$ inside one single parent relativistic model, which can actually also exactly include Fendley-Intriligator, a.k.a. Sine-Gordon at a special value, showing the common root in terms of fractional statistics. We cannot but entertain the hope that our interpolations are perhaps related to the ones which are being pursued in the literature from the worldsheet viewpoint, such as [4, 31, 32, 29] and the very resembling [17], while we are here focused on trying to follow the tracks left by the simulacrum of the Sine-Gordon $S$-matrix. This paper only treats one block (left-left) for lightness, as if we are displaying a principle and not a complete armamentary, but for completeness one should extend to all chirality blocks - deriving an interpolating crossing equation and parity properties will be particularly sensitive to that assemblage. Incorporating bound states [23] will also be important.
@@
It would also be very interesting to ascertain whether there is a natural way to identify the fractional statistics of the excitations within the structure of the QSC [18, 19, 20].
2 Sine-Gordon at special coupling and $AdS_{3}$
@@
Let us begin by demonstrating how the Sine-Gordon model taken at a special value of the coupling develops an ${\cal{N}}=2$ supersymmetry. More precisely, the statement which we will prove - and which was first observed by [47] - is the following: at the special coupling $\beta^{3}=\frac{16\pi}{3}$ the usual quantum group symmetry of the Sine-Gordon model becomes isomorphic to an ${\cal{N}}=2$ supersymmetry in the soliton-antisoliton sector [^3] As we will shortly review, this is completely internal to Sine-Gordon itself, spontaneously emergent so to say. This is not what is often called ${\cal{N}}=2$ supersymmetric Sine-Gordon (references can be found for instance in [14]), which is instead an extension of Sine-Gordon including more fields to make the Lagrangian supersymmetric to start with. In our case, the Lagrangian is the good old Sine-Gordon, and the supersymmetry happens as an accident for the special value of $\beta$., with fractional fermionic number assignment of the states $|\phi|=\mbox{soliton}=\frac{1}{2}$ and $|\psi|=\mbox{antisoliton}=-\frac{1}{2}$.
@@
Let us recall that the soliton-antisoliton Sine-Gordon $S$-matrix (for a review see for instance [51]) takes the form
$S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&S_{R}&S_{T}&0\\ 0&S_{T}&S_{R}&0\\ 0&0&0&1\end{pmatrix}$, (2.1)
@@
$S_{T}=\frac{\sinh\frac{\pi\theta}{\xi}}{\sinh\frac{\pi(i\pi-\theta)}{\xi}},\qquad S_{R}=\frac{i\sin\frac{\pi^{2}}{\xi}}{\sinh\frac{\pi(i\pi-\theta)}{\xi}},\qquad\theta=\theta_{1}-\theta_{2},$ (2.2)
@@
$|\phi\rangle=\begin{pmatrix}1\\ 0\end{pmatrix},\qquad|\psi\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$, (2.3)
@@
$\Delta_{21}\hat{E}\,\,S=S\,\,\Delta\hat{E}$, (2.4)
with
$\Delta\hat{E}=\hat{E}\otimes\mathbbmss{1}+q^{-2H}\otimes\hat{E}$, (2.5)
and where $\hat{E}$ is represented by the matrix $x_{i}E_{\phi\psi}$, having set $x_{i}=\exp[\frac{\pi\theta_{i}}{\xi}]$, for the particle $i=1,2$. $E_{\rho\tau}$ is the matrix with only one non-zero entry equal to $1$ in row $\rho$, column $\tau$. We also have $q=\exp\Big{[}\frac{8\pi^{2}i}{\beta^{2}}\Big{]}$, and $H=\frac{1}{2}(E_{\phi\phi}-E_{\psi\psi})$.
@@
$S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$. (2.6)
@@
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(\hat{E}\otimes\mathbbmss{1}+q^{-2H}\otimes\hat{E})|\phi\rangle\otimes|\psi\rangle=0+(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle,$ (2.7)
@@
Both $|\phi\rangle$ and $|\psi\rangle$ have the same fermionic number $|\phi|=|\psi|=0$ in Sine-Gordon, and therefore $\hat{E}$ has zero fermionic number - because it sends the state $|\psi\rangle$ into the state $|\phi\rangle$ [^4] The fact that soliton and anti-soliton Bethe eigenstates behave like fermions, even though the bare statistics of the vectors in this original picture is 0, is due the the Sine-Gordon $S$-matrix satisfying $S_{\rho\rho}(0)=-1$, $\rho=\phi,\psi$ [52]..
We can now rephrase (2.7) in a new (fractional) picture:
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\phi\rangle\otimes|\psi\rangle=(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle+0.$ (2.8)
@@
Expressions (2.7) and (2.8) giving the same numerical vector is achieved by this: first, $\hat{E}$ is represented as $x_{i}E_{\phi\psi}=x_{i}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$, but has fermionic number $|\hat{E}|=1$; second, $|\phi\rangle$ is represented as $\begin{pmatrix}1\\ 0\end{pmatrix}$, but has fermionic number $|\phi|=\frac{1}{2}$; third, $|\psi\rangle$ is represented as $\begin{pmatrix}0\\ 1\end{pmatrix}$, but has fermionic number $|\psi|=-\frac{1}{2}$. This means that
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle=(-)^{|\hat{E}||1|}|\phi\rangle\otimes\hat{E}|\psi\rangle,$ (2.9)
which gives the same result.
@@
This exercise can be repeated for a partner supercharge $\hat{F}=x_{i}\,E_{\psi\phi}$, $\Delta\hat{F}=\hat{F}\otimes\mathbbmss{1}+q^{2H}\otimes\hat{F}$. We shall not need the entire ${\cal{N}}=2$ algebra for our scopes, but just half will suffice. The extension of this argument to all the other generators and all other states follows. This reveals the emergence of supersymmetry at this special value of the coupling: we have managed to rephrase the action of the coproduct of the quantum group generators into an equivalent form involving a supercharge, which consistently acts on states with fermionic number $\pm\frac{1}{2}$ in the (graded) tensor product. The coproduct action has become local - as exemplified by (2.8).
@@
In some sense all that has happened is that the quantum-group deformation $q^{-2H}$, resp. $q^{2H}$, in the coproduct at this value of the coupling become equal to $(-1)^{|\hat{E}|\begin{pmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{pmatrix}}$, resp. $(-1)^{|\hat{F}|\begin{pmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{pmatrix}}$, which simulate the statistical factors which a fractional supersymmetry would produce when swapping the corresponding generators with the states. We have again opted for using the same symbol for the generators represented by the same matrices, despite the fact that they might have different fermionic grading, again hoping that the distinction will always be clear and to lighten the notation. This way it will always be
$\hat{E}\propto E_{\phi\psi},\qquad\hat{F}\propto E_{\psi\phi}.$ (2.10)
2.1 Relation with massless pure-RR $AdS_{3}\times S^{3}\times T^{4}$

### v4

@@ Abstract
The massless $S$-matrix of the pure RR $AdS_{3}\times S^{3}\times T^{4}$ theory is very similar but also crucially distinct from Fendley-Intriligator’s famous ${\cal{N}}=2$ $S$-matrix, in turn related to Sine-Gordon taken at a special coupling. In this article we review the reason why supersymmetry emerges but with a fractional statistics of the particles. We then use this to obtain an interpolating $S$-matrix between massless $AdS_{3}$ and Sine-Gordon, and further find two interpolating $S$-matrices between these and the mixed-flux relativistic $S$-matrix, solving the Yang-Baxter equation all the way in-between. They are all relativistic invariant and braiding unitary, can be written in terms of particles with fractional statistics, and embed all the $AdS_{3}$ models into a parent integrable system. This also achieves the reformulation of all the $AdS_{3}$ integrable scattering problems in the domain of fractional statistics. [^1] DMUS-MP-25/04.
I Introduction and Conclusions
It is undeniable that integrability of strings propagating on $AdS_{3}\times S^{3}\times T^{4}$ [1] (for reviews see [2, 3, 4]) continues to offer opportunities for surprises, keeping on with the tradition inaugurated by $AdS_{5}$ and $AdS_{4}$ [5]. The exact $S$-matrix theory is one of the richest available [6, 7, 8, 9, 10, 11] and the presence of massless modes [12, 13, 14], see also [15] and [16, 17], keeps pushing the technology further and further. The Quantum Spectral Curve (QSC) [18], see also [19], is undergoing an expansion of its methods to accommodate this model [20]. Meanwhile [21], see also [22, 23, 24], have considerably revitalised the field with new algebraic data and have created an exciting race with significant progress on all sides [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
@@ I Introduction and Conclusions
$E=\sqrt{\Big(m+\frac{k}{2\pi}p\Big)^{2}+4h^{2}\sin^{2}\frac{p}{2}}.$ (1)
@@
Notably $k\in\mathbbmss${Z} stands for the WZW level and $m$ is for the mass of the particle. An important wealth of work on this and similar deformations has now been done [43]. As it is customary, one calls pure RR the $k=0$ case (and mixed-flux otherwise).
@@
There is a special subarea inside this field currently involving a very intense activity, and that is the relativistic sector. This goes from the BMN limit being non-trivial for same-chirality massless scattering [14], where it describes a CFT á lá Zamolodchikov [44], to the entire massless sector being of difference form [35, 36, 13, 21], to the mixed-flux scattering being amenable to a variety of massive relativistic limits [36, 23, 24], all pointing toward famous integrable $S$-matrices of the past - with crucial differences. The observation [14] is that the massless $S$-matrix of the pure RR theory is very similar but not quite the same as Fendley-Intriligator’s ${\cal{N}}=2$ $S$-matrix [45, 46, 47]. In turn, this ${\cal{N}}=2$ $S$-matrix is related to Sine-Gordon taken at a special coupling [48], and it explicitly appeared in the paper by Mussardo [49], where important properties in connection with Sine-Gordon at a special coupling are outlined. Considerations on the connections with the ${\cal{N}}=2$ literature in the case of mixed-flux are also expanded upon in [23], and in particular in [24], where the emphasis is on the differences in the global structure of the supersymmetry algebra.
@@
The focus of this paper is instead the difference in the nature itself of the supersymmetry, in particular the appearance of fractional statistics. We review the detailed reason why supersymmetry emerges but is not exactly that which one has in pure RR $AdS_{3}$, due to a different assignment of fermionic number of the particles. We use the coproduct of the underlying quantum affine symmetry of Sine-Gordon to show this, following the presentation in Bernard and LeClair [48], see also section 2 of [45] and [50]. We use this to obtain an interpolating $S$-matrix between massless $AdS_{3}$ and Sine-Gordon.
@@
Then we deformed this interpolating $S$-matrix in order to relax the special coupling and let it be generic. The aim was to obtain an interpolation to the (or one of the) mixed flux $S$-matrix (-ces). We managed to find two distinct classes of interpolations which solve the Yang-Baxter equation (YBE) all the way in-between. The first class, denoted by $S_{\alpha}$, corresponds to deformations of Sine-Gordon at generic $q$, and contains the Fendley-Intriligator case and the massless $AdS_{3}$ as limiting cases. The second class, denoted by $S_{\alpha+}$ in its most general incarnation, corresponds to deformations of Sine-Gordon at the special coupling q=iq=-i, and contains the Fendley-Intriligator case, the massless $AdS_{3}$ and the relativistic mixed-flux as limiting cases. They are all relativistic invariant, and satisfy phyical unitarity for a choice of parameters. Both classes display fractional statistics for the scattering particles. These $S$-matrices are all regular, namely $S(\theta)\to\mathbbmss{1}\otimes\mathbbmss{1}$ as $\theta\to 0$, , with $\theta=\theta_{1}-\theta_{2}$ being the relativistic rapidity-difference of the two scattering particles. This means that they sit within the classification of [51], possibly section 4.3.1 modulo twists  [^2] They are also all of free-fermion type, which is consistent with having all the AdS/CFT $S$-matrix of free-fermion type - and also the Sine-Gordon $S$-matrix at special coupling $\xi\to 2\pi$, since Sine-Gordon satisfies $\mbox{const.}=4\cos^{2}\frac{\pi^{2}}{\xi}$ in formula (2.2) of [31]. They all have $\mathfrak{sl}_{2}$-type quantum symmetry, they are braiding unitarity in deference to their quantum group origin, and they can be rephrased in terms of particles with fractional statistics  [^3] Statistics in $1+1$ dimension is a fascinating subject, see for instance [52, 34].. Inspired by [34], we highlight this fact by calling them all fractional $S$-matrices - what seems to be a ubiquitous feature at this point [46]. We believe that this work serves a useful purpose to explicitly embed all the relativistic $S$-matrices appearing from the study of string theory in $AdS_{3}$ inside one single parent relativistic model, which can actually also exactly include Fendley-Intriligator, a.k.a. Sine-Gordon at a special value, showing the common root in terms of fractional statistics. It will be of paramount importance to relate our interpolations with the ones which are being pursued in the literature from the worldsheet viewpoint, such as [4, 32, 33, 29] and the very resembling [17]. This paper treats one particular block, corresponding to the left-moving kinematics for both scattering particles - in Zamolodchikov’s picture of massless scattering [44], this is a non-perturbative effect which does not correspond to a perturbative classical configuration. We choose to explicitly display one particular block for definiteness and since we are describing a principle. Incorporating different blocks and bound states [23] shall promptly follow.
@@
It would also be very interesting to ascertain whether there is a natural way to identify the fractional statistics of the excitations within the structure of the QSC as constructed in [18, 19, 20].
II Sine-Gordon at special coupling and $AdS_{3}$
@@
Let us begin by demonstrating how the Sine-Gordon model taken at a special value of the coupling develops an ${\cal{N}}=2$ supersymmetry. More precisely, the statement which we will review the proof of is the following: at the special coupling $\beta^{2}=\frac{16\pi}{3}$ the usual quantum group symmetry of the Sine-Gordon model becomes isomorphic to an ${\cal{N}}=2$ supersymmetry in the soliton-antisoliton sector  [^4] As we will shortly review, this is completely internal to Sine-Gordon itself, spontaneously emergent so to say. This is not what is often called ${\cal{N}}=2$ supersymmetric Sine-Gordon (references can be found for instance in [14]), which is instead an extension of Sine-Gordon including more fields to make the Lagrangian supersymmetric to start with. In our case, the Lagrangian is exclusively Sine-Gordon, and the supersymmetry emerges for the special value of $\beta$., with fractional fermionic number assignment of the states $|\phi|=\mbox{soliton}=\frac{1}{2}$ and $|\psi|=\mbox{antisoliton}=-\frac{1}{2}$.
@@
Let us recall that the soliton-antisoliton Sine-Gordon $S$-matrix [53] (for a review see for instance [54]) takes the form
$S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&S_{R}&S_{T}&0\\ 0&S_{T}&S_{R}&0\\ 0&0&0&1\end{pmatrix}$, (2)
@@
$S_{T}=\frac{\sinh\frac{\pi\theta}{\xi}}{\sinh\frac{\pi(i\pi-\theta)}{\xi}},\qquad S_{R}=\frac{i\sin\frac{\pi^{2}}{\xi}}{\sinh\frac{\pi(i\pi-\theta)}{\xi}},$
$\qquad\theta=\theta_{1}-\theta_{2},$ (3)
@@
$|\phi\rangle=\begin{pmatrix}1\\ 0\end{pmatrix},\qquad|\psi\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$, (4)
@@
$\Delta_{21}\hat{E}\,\,S=S\,\,\Delta\hat{E}$, (5)
where $\Delta_{21}$ is the coproduct applied to the permuted combination of particles $2\otimes 1$, and with
$\Delta\hat{E}=\hat{E}\otimes\mathbbmss{1}+q^{-2H}\otimes\hat{E}$, (6)
and where $\hat{E}$ is represented by the matrix $x_{i}E_{\phi\psi}$, having set $x_{i}=\exp[\frac{\pi\theta_{i}}{\xi}]$, for the particle $i=1,2$. $E_{\rho\tau}$ is the matrix with only one non-zero entry equal to $1$ in row $\rho$, column $\tau$. We also have $q=\exp[\frac{8\pi^{2}i}{\beta^{2}}\Big]$, and $H=\frac{1}{2}(E_{\phi\phi}-E_{\psi\psi})$. We then recall that $\Delta(.)$ is the coproduct, in such a way that (5) is our notation for the standard condition of invariance of the $S$-matrix under the two-particle representation of the symmetry [48].
@@
$S=\Phi(\theta)\begin{pmatrix}1&0&0&0\\ 0&\mbox{sech}\frac{\theta}{2}&-i\tanh\frac{\theta}{2}&0\\ 0&-i\tanh\frac{\theta}{2}&\mbox{sech}\frac{\theta}{2}&0\\ 0&0&0&1\end{pmatrix}$. (7)
@@
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(\hat{E}\otimes\mathbbmss{1}+q^{-2H}\otimes\hat{E})|\phi\rangle\otimes|\psi\rangle$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt=0+(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle,$ (8)
@@
Both $|\phi\rangle$ and $|\psi\rangle$ have the same fermionic number $|\phi|=|\psi|=0$ in Sine-Gordon, and therefore $\hat{E}$ has zero fermionic number - because it sends the state $|\psi\rangle$ into the state $|\phi\rangle$  [^5] The fact that soliton and anti-soliton Bethe eigenstates behave like fermions, even though the bare statistics of the vectors in this original picture is 0, is due the Sine-Gordon $S$-matrix satisfying $S_{\rho\rho}(0)=-1$, $\rho=\phi,\psi$ [55]..
We can now rephrase (8) in a new (fractional) picture:
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(\mathbbmss{1}\otimes\hat{E}+\hat{E}\otimes\mathbbmss{1})|\phi\rangle\otimes|\psi\rangle$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt=(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle+0.$ (9)
@@
Expressions (8) and (9) giving the same numerical vector is achieved by this: first, $\hat{E}$ is represented as $x_{i}E_{\phi\psi}=x_{i}\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$, but has fermionic number $|\hat{E}|=1$; second, $|\phi\rangle$ is represented as $\begin{pmatrix}1\\ 0\end{pmatrix}$, but has fermionic number $|\phi|=\frac{1}{2}$; third, $|\psi\rangle$ is represented as $\begin{pmatrix}0\\ 1\end{pmatrix}$, but has fermionic number $|\psi|=-\frac{1}{2}$. This means that
$\Delta\hat{E}|\phi\rangle\otimes|\psi\rangle=(-1)^{\frac{1}{2}}\,x_{2}\,|\phi\rangle\otimes|\phi\rangle$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt=(-)^{|\hat{E}||1|}|\phi\rangle\otimes\hat{E}|\psi\rangle,$ (10)
which gives the same result. We see that we have been able to equivalently rephrase the contribution from $q^{-2H}$ in the original coproduct, now in terms of a fractional fermionic number, carried by a redefined value for $|1|$, that is $|1|=|\phi|=\frac{1}{2}$.
@@
This exercise can be repeated for a partner supercharge $\hat{F}=x_{i}\,E_{\psi\phi}$, $\Delta\hat{F}=\hat{F}\otimes\mathbbmss{1}+q^{2H}\otimes\hat{F}$. We shall not need the entire ${\cal{N}}=2$ algebra for our scope, but just half will suffice. The extension of this argument to all the other generators and all other states follows. This reveals the emergence of supersymmetry at this special value of the coupling: we have managed to rephrase the action of the coproduct of the quantum group generators into an equivalent form involving a supercharge, which consistently acts on states with fermionic number $\pm\frac{1}{2}$ in the (graded) tensor product. The coproduct action has become local - as exemplified by (9).
@@
In some sense all that has happened is that the quantum-group deformation $q^{-2H}$, resp. $q^{2H}$, in the coproduct at this value of the coupling become equal to $(-1)^{|\hat{E}|H}$, resp. $(-1)^{|\hat{F}|H}$, with $H=\begin{pmatrix}\frac{1}{2}&0\\ 0&-\frac{1}{2}\end{pmatrix}$, which simulate the statistical factors which a fractional supersymmetry would produce when swapping the corresponding generators with the states. We have again opted for using the same symbol for the generators represented by the same matrices, despite the fact that they might have different fermionic grading, again hoping that the distinction will always be clear and to lighten the notation. This way it will always be
$\hat{E}\propto E_{\phi\psi},\hskip 18.49988pt\hat{F}\propto E_{\psi\phi}.$ (11)
II.1 Relation with massless pure-RR $AdS_{3}\times S^{3}\times T^{4}$